OpenStudy (crashonce):

A, 2, B are in AP and A, 5, B are in GP. The quadratic equation whose roots are A and B is?

OpenStudy (loser66):

What do AP and GP stand for?

OpenStudy (crashonce):

artithmetic progression and geometric progression

OpenStudy (mathmath333):

\[\frac{ A+B }{ 2 }=2\\A\times B=5^2\]

OpenStudy (triciaal):

for a quadratic equation ax^2 + bx + c = 0 the product of the roots = ac and the sum of the roots = b given the sum of the roots A and B = 4 and the product of the roots, AB = 25 substitute in the equation

OpenStudy (mathmath333):

u need to do \((x-a)(x-b)=0\\ x^2-x(a+b)+ab=0\)